OpenStudy (anonymous):

multiply: x^3/4 times x^1/6

Parth (parthkohli):

Multiply the numerators and denominators.\[\large{x^3 \over 4} \times {x \over 6} \implies {x^3 \times x \over 4 \times 6} \]Does that help you?

OpenStudy (ghazi):

nope @ParthKohli

OpenStudy (ghazi):

x^(3/4+1/6)

Parth (parthkohli):

Oh, the question has exponents.

OpenStudy (ghazi):

x^11/12 i guess

Parth (parthkohli):

\[x^{6} \times {x^{4}} \implies x^{6 + 4} \implies {x^{10}} \]Does this example help?

OpenStudy (anonymous):

@ghazi YESS

OpenStudy (ghazi):

:)

Parth (parthkohli):

Oh... no... no... noo....

OpenStudy (ghazi):

@ParthKohli

OpenStudy (ghazi):

?? what has happened to you?

Parth (parthkohli):

:(